If josh is 4 and his younger bother is half that
How old is his little brother ?
And if josh is 100 how old is his little brother?

Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

If josh is 4 and his younger bother is half that
How old is his little brother ?
4/2 = 2
---------------
And if josh is 100 how old is his little brother?
If it's the same pair, 100 - 2 = 98
----------
If it's a separate problem, 100/2 = 50

Answer by addingup(3677)

You need to learn to either divide by 2 or multiply times 0.5 or 1/2